OpenStudy (anonymous):

is the answer pi/4 Find the period of the function. y = -3 cos 1/4x pi/4 -3 3pi/4 8π

random231 (random231):

nope. think again!

OpenStudy (anonymous):

oh is it -3?? i got amplitude and period mixed up because amp is always positive

OpenStudy (anonymous):

You need to write down the form, of how a cossine function looks. It would be: y= A * cos(w*x), where A is the amplitude and w is the angular frequency, w= 2*pi/ T, where T is the period of the function. So by looking at it you can tell that 1/4 = 2*pi/T.

random231 (random231):

@Barsukis is right!